Use r=2√(5d).
A car's rate (r) is 55mi/h. How far (d) will it skid?

Answer by stanbon(75887)

Use r=2√(5d).
A car's rate (r) is 55mi/h. How far (d) will it skid?
--------
55 = 2*sqrt(5d)
sqrt(5d) = (55/2)
---
5d = (55/2)^2
---
d = (11*55)/4
---
d = 151.25 ft
==========================
Cheers,
Stan H.
----------------------